# Cascade-driven opinion dynamics on social networks

### v1

@@ Abstract
Online social networks (OSNs) have transformed the way individuals fulfill their social needs and consume information. As OSNs become increasingly prominent sources for news dissemination, individuals often encounter content that influences their opinions through both direct interactions and broader network dynamics. In this paper, we propose the Friedkin-Johnsen on Cascade (FJC) model, which is, to the best of our knowledge, is the first attempt to integrate information cascades and opinion dynamics, specifically using the very popular Friedkin-Johnsen model. Our model, validated over real social cascades, highlights how the convergence of socialization and sharing news on these platforms can disrupt opinion evolution dynamics typically observed in offline settings. Our findings demonstrate that these cascades can amplify the influence of central opinion leaders, making them more resistant to divergent viewpoints, even when challenged by a critical mass of dissenting opinions. This research underscores the importance of understanding the interplay between social dynamics and information flow in shaping public discourse in the digital age.
@@ I Introduction
Socialization represents a fundamental aspect of human behavior, with empirical evidence consistently demonstrating the correlation between human health and social engagements [1, 2, 3], underscoring how isolation has serious consequences on both physical and mental health conditions [4, 5, 6], and could potentially lead to premature mortality [7]. The advent of online social networks (OSNs) has significantly reshaped how individuals fulfill their social needs, offering unprecedented opportunities to connect with others. Additionally, OSNs have emerged as prominent sources of real-time news dissemination, with a substantial proportion of the population relying on these platforms for information [8]. Political figures, journalists, and ordinary users frequently utilize OSNs to disseminate announcements and opinions, rendering them integral to contemporary information consumption patterns. Moreover, media organizations have implemented strategies to encourage the sharing of their content on OSNs, facilitating the seamless integration of news into users’ social interactions. Therefore, during our social activities online, we can also be exposed to the news even incidentally. However, this convergence of socialization and news consumption on OSNs has raised concerns regarding the formation of filter bubbles and echo chambers, wherein individuals are exposed primarily to content that reinforces their existing beliefs. Such environments pose risks of fostering biased perspectives and perpetuating misinformation, particularly with the proliferation of fake news campaigns designed to manipulate public opinion [9]. These dynamics underscore the complex interplay between socialization, news consumption, and information dissemination in the digital era, warranting further empirical investigation and theoretical exploration within interdisciplinary frameworks.
@@
The dynamics of opinion formation represent a multifaceted research area situated at the intersection of various disciplines, including sociology, economics, psychology, control theory, computational sociology, and physics. Numerous models have been proposed to describe the mechanisms underlying opinion dynamics, aiming to capture observed theoretical and empirical phenomena. These models often account for factors such as individuals’ inclination to reduce cognitive dissonance with close acquaintances or their resistance to relinquishing deeply held biases. Among the prominent models in this domain is the Friedkin-Johnsen (FJ) model [10] belonging to the category of averaging models [11, 12, 10], according to which people iteratively adjust their opinions towards the average opinion of their neighbours, with the strength of social ties weighting the influence of each neighbor’s opinion. Additionally, the FJ model incorporates a parameter representing individuals’ open-mindedness, referred to as susceptibility, which modulates the extent to which individuals revise their opinions. The FJ model has gained popularity due to its mathematical tractability and its empirical validation on real social networks. It has demonstrated the ability to replicate various observed mechanisms in social networks, including the phenomenon of wisdom of the crowd [13] and the emergence of polarization [14] and recently, it has also been proposed as the building block to model multilateral international negotiation processes [15].
The FJ model has undergone rigorous validation in offline scenarios, where opinions evolve reflecting real-world interactions and debates on various topics. However, the dynamics change when socializing online, as we are exposed to friends’ opinions not through direct interaction but via their public activity on social media platforms, such as posting or commenting on news articles. Consequently, we may frequently encounter cognitive dissonance toward opinions that are expressed more frequently online. These online opinions often garner larger-scale attention and spread more intensely than their offline counterparts, potentially heightening the impact of social opinion exchange in digital social environments. These opinions are conveyed via textual posts, resharing of news items, or images. The dynamics of how content items are shared and spread online are studied under the name of information cascades [16].
@@
Opinion dynamics and information cascades operate at distinct granularity and temporal scales. Information cascades focus on how individual pieces of information (e.g., a post, an image) are shared and propagate within the social network, while opinion dynamics concern the evolution of opinions through interactions among individuals. Both processes exhibit emergent behaviors, transitioning from local interactions to global effects such as content virality and polarization. In opinion dynamics, the specific social interactions leading to opinion change remain abstract and unspecified. Conversely, information cascades emphasize the sharing mechanism of content without addressing the effects of such sharing. However, given that reposting online content received from friends is a common online interaction, it is reasonable to expect an intertwining of opinion dynamics and information cascades. Despite this, the two processes have thus far been treated separately in the literature. In this work, we set out to fill this gap and establish an integrated model whereby the microscopic act of content sharing bears an effect in the macroscopic phenomenon of opinion evolution in online social networks.
To address this issue, we propose a novel model that integrates the FJ model with information cascades. Specifically, we hypothesize that opinions on a selected topic are expressed following the dissemination flow of online posts. Initially shared by certain sources, the post propagates as a cascade to nodes within their immediate and subsequent neighborhoods until it loses interest or reaches all nodes. Each shared post is assumed to convey the opinion of the node on the selected topic, e.g., expressed in text as feedback to a shared news article, thereby triggering cognitive responses from neighbors. The reception of a post from a neighbor, triggering an opinion update on the receiving node.
Furthermore, building upon our previous work [14], we analyze the impact of the cascade on polarization, evaluating the polarization values on the opinion vectors that maximize polarization of the classical FJ model. We tested the model over synthetic and real social networks. Our research reveals that the cascade-driven FJ model amplifies the influence of network topology compared to the classical FJ model. This empowers central nodes to act as opinion leaders, potentially amplifying or mitigating network polarization. This result implies that sharing mechanisms on online social networks can introduce imbalances in the opinion dynamics process compared to offline socialization, possibly resulting in higher levels of network polarization. These findings underscore the potential for online social networks to drive change and shape our social interactions. Finally, we validated our model with a dataset of real cascade on the Twitter platform and proved that the model is able to predict very well the real opinion dynamics. With our model, we can better understand the dynamics of online social networks and leverage this knowledge to create more effective communication strategies for businesses, organizations, and individuals alike.
@@ II Related Works
The modeling of opinion dynamics has garnered the interest of researchers for decades, with this interest further amplifying with the emergence of online social networks. There are several classes of opinion dynamics models, and the literature on the topic is extensive [17]. Here, we concentrate on the class of averaging models, which posit that users adjust their opinions based on those of the people around them to mitigate social discomfort. Among these models, the FJ model [10] is one of the most studied [18, 19, 20, 21, 22, 23, 24, 25, 26, 14]. Its inner workings are explained in Seciton III-C. Recently, research on the FJ model has taken two main directions. The first involves optimization problems based on FJ [27, 28, 29, 21, 18, 19], whose main goal is to make interventions into the social network to achieve specific desired outcomes in the final opinions. For example, some investigate how to modify the initial opinion of a set of nodes to maximise the opinion discord [27] or minimize the overall opinion [28]. The second direction explores the possible extensions of the FJ model to capture diverse observed features in the online social networks [30, 31, 32, 33, 34]. In particular, these papers incorporate both attractive and repulsive relationships between users [31, 33], even assuming that the influence of a node on another can vary according to the homophily principle [34]. Furthermore, Bhalla et al. in [30] analyzes the polarization effect of FJ on time-varying networks and Chitra et al. in [32] extended the FJ to reproduce filter bubbles. Differently from these works, we want to investigate the effect of the classical FJ model in a static network when altering the social debate, which is regulated by the posting flow of news on online social networks.
@@
To model the posting flow, we mainly take hints from the very popular independent-cascade model in [35] (see Section III-B for a description of how it works), which is able to reproduce realistic diffusion dynamics [36]. This seminal work has inspired many following studies and extentions [37, 38, 39, 40], which extend the classical model to include collaborative and/or antagonistic relationships among users or dynamic networks. The main goal of information diffusion is to study the rate of diffusion of viral contents depending on the parameters. Instead, we are not interested in the popularity of contents, but rather in their diffusion process. Hence, we used the original independent-cascade model and analyzed its impact on the opinion formation process. Similarly to opinion dynamics, the literature on information diffusion is extensive, so we refer the interested reader to dedicated surveys for more details [41].
The work closer to ours is [42] which proposes a model to estimate the inter-dependencies of viral contents and opinion formation. This model operates with two mechanisms during each time round.The first mechanism involves opinion updating, modeled using a relaxed version of the FJ model, which cannot reproduce polarization (for further details, see [14]). The second one is the proper interplay between information spreading and opinion adjustments. In this mechanism, the content spreads and its resulting popularity acts as a probability for nodes to change their innate opinion. In this model, the information cascade is not linked to people’s opinions when resharing the contents, nor does the flow of content sharing affect opinion evolution. In contrast, we assume that sharing content reflects peoples’ opinions and that the sharing process itself impacts opinion updating. Our goal is to understand how the propagation flow alters the FJ and investigate the impact on polarization. As far as we know, our work is the first approach to investigate this problem.
@@ III-A The social network
We can describe an online social network platform as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the $v$ nodes $\mathcal{V}$ represent users and the edges $\mathcal{E}$ represent social connections between them. In the most general form, the graph $G$ is directed and the edges $(i,j)$ represent a follow relationships. Specifically, $(i,j)\in\mathcal{E}$ means that $i$ follows $j$, thus $j$ influences $i$ as its activity appears in $i$’s news feed. The set of nodes that has an influence on $i$ is thus composed of $i$’s out-neighbours, which we will denote with $\mathcal{N}^{+}(i)$. Instead, the set of nodes that is influenced by node $j$ is the in-neighbourhood of $j$, which we indicate with $\mathcal{N}^{-}(i)$ [^1] Please note that, generally, the neighbourhood of a node corresponds to our out-neighbourhood.. It is worth noting that in the special case of an undirected node, $\mathcal{N}^{+}(i)$ and $\mathcal{N}^{-}(i)$ are the same. This social connection can arise, for example, when $i$ follows $j$ on platforms like X or Facebook or when they are friends on Facebook. The strength of this relationship is mathematically represented by the quantity $\hat{w}_{ij}$, which is the $(i,j)$-th entry of the $n\times n$ matrix $\hat{W}$ that we call social matrix. The weight that node $i$ gives to its out-neighbours represents the degree of influence of their opinions. It is provided by the $(i,j)$-th entry of the $n\times n$ matrix $W$, named the influence matrix $W$, which is row-stochastic. On the $i$-th row, for example, we can see how the influence on $i$’s opinion is distributed among all nodes in $i$’s network. In situations where additional data is lacking, the matrix $W$ can be derived from $\hat{W}$ through row-normalization, expressed as $w_{ij}=\frac{\hat{w}{ij}}{\sum_{j}\hat{w}{ij}}$. This process capitalizes on the insight that stronger relationships should carry greater weight in opinion formation. It’s worth noting that, with this mechanism, even in cases where the social graph exhibits symmetry, the influence matrix results asymmetric.
@@ III-B Information cascade
Our model conceptualizes the act of posting social content following the independent cascade model [35]. In the independent cascade model, each node can be active or inactive. An initial set of active nodes $v$ have a single chance to activate their neighbours $w$, and each activation is successful with probability $p_{vw}$. The nodes $w$ that are successfully activated, become active at the next step and try to activate their neighbours as well. The process continues until no other nodes can be activated.
@@ III-C Friedkin-Johnsen opinion dynamic model
The FJ opinion dynamic model is a mathematical model that describes how people’s opinions change over time due to their mutual influences. It does this by representing people’s opinions as real numbers and then modelling two opposing forces that affect how those opinions evolve. The first force is an initial bias that each person $i$ has, which is represented by $u_{i}$. The second force is a desire to reduce the discomfort people feel when their opinions differ from those around them, as described by the cognitive dissonance theory. Each person $i$ has a parameter $\lambda_{i}$, called susceptibilty, that determines how much weight they give to each of these forces. If $\lambda_{i}$ is 0, it means the person is completely stubborn and will never change its mind. If $\lambda_{i}$ is 1, it means the person is completely naive and will totally get over their initial bias. Node susceptibilities are generally represented by the diagonal matrix $\Lambda$, whose $i$-th diagonal element is equal to the susceptibility $\lambda_{i}$.
The FJ model is mathematically expressed as a simple update equation that is applied to each person’s opinion at discrete times. As discussed in our previous work [14], in the literature, the term “FJ model” refers to many variants of the original model. The update expression for the most general variant of FJ model is given by the following equation:
@@
while the rest of the nodes maintain their opinion. Here matrices $H=\{h_{ij}\}$ and $\Gamma=\{\gamma_{ij}\}$ replace $\Lambda$ and $W$ of the synchronous model, and are different from them. In fact, to guarantee the convergence of the model in Equation (3) to that in Equation (1), the formulation has to take into account the probability of node $i$ being randomly selected, which, in turn, depends on the degree of nodes. Since high-degree nodes are involved in more edges, when selecting edges at random high-degree nodes have a higher probability of being selected. For this reason, $H$ and $\Gamma$ should be defined from $\Lambda$ and $W$ as follows:

### v2

@@ Abstract
Online social networks (OSNs) have transformed the way individuals fulfill their social needs and consume information. As OSNs become increasingly prominent sources for news dissemination, individuals often encounter content that influences their opinions through both direct interactions and broader network dynamics. In this paper, we propose the Friedkin-Johnsen on Cascade (FJC) model, which, to the best of our knowledge, is the first attempt to integrate information cascades and opinion dynamics, specifically using the very popular Friedkin-Johnsen model. Our model, validated over real social cascades, highlights how the convergence of socialization and sharing news on these platforms can disrupt opinion evolution dynamics typically observed in offline settings. Our findings demonstrate that these cascades can amplify the influence of central opinion leaders, making them more resistant to divergent viewpoints, even when challenged by a critical mass of dissenting opinions. This research underscores the importance of understanding the interplay between social dynamics and information flow in shaping public discourse in the digital age.
@@ I Introduction
Online social networks (OSNs) have fundamentally transformed how individuals communicate, share information, and form opinions in the digital age. Serving both as platforms for social interaction and as primary sources of real-time news, OSNs facilitate unprecedented levels of connectivity and information dissemination [1]. Additionally, OSNs have emerged as prominent sources of real-time news dissemination, with a substantial proportion of the population relying on these platforms for information [1]. Political figures, journalists, and ordinary users frequently utilize OSNs to disseminate announcements and opinions, rendering them integral to contemporary information consumption patterns. Moreover, media organizations have implemented strategies to encourage the sharing of their content on OSNs, facilitating the seamless integration of news into users’ social interactions. Therefore, during our social activities online, we can also be exposed to the news even incidentally. However, this convergence of socialization and news consumption on OSNs has raised concerns regarding the formation of filter bubbles and echo chambers, wherein individuals are exposed primarily to content that reinforces their existing beliefs. Such environments pose risks of fostering biased perspectives and perpetuating misinformation, particularly with the proliferation of fake news campaigns designed to manipulate public opinion [2]. These dynamics underscore the complex interplay between socialization, news consumption, and information dissemination in the digital era, warranting further empirical investigation and theoretical exploration within interdisciplinary frameworks. Understanding the mechanisms behind these phenomena requires an integrated study of how opinions evolve through social influence and how information cascades propagate within these networks.
@@
The dynamics of opinion formation represent a multifaceted research area situated at the intersection of various disciplines, including sociology, economics, psychology, control theory, computer science, computational social science, and physics. Numerous models have been proposed to describe the mechanisms underlying opinion dynamics, aiming to capture observed theoretical and empirical phenomena. These models often account for factors such as individuals’ inclination to reduce cognitive dissonance with close acquaintances or their resistance to relinquishing deeply held biases. Among the prominent models in this domain is the Friedkin-Johnsen (FJ) model [3] belonging to the category of averaging models [4, 5, 3], according to which people iteratively adjust their opinions towards the average opinion of their neighbours, with the strength of social ties weighting the influence of each neighbor’s opinion. Additionally, the FJ model incorporates a parameter representing individuals’ open-mindedness, referred to as susceptibility, which modulates the extent to which individuals revise their opinions. The FJ model has gained popularity due to its mathematical tractability and its empirical validation on real social networks. It has demonstrated the ability to replicate various observed mechanisms in social networks, including the phenomenon of wisdom of the crowd [6] and the emergence of polarization [7] and recently, it has also been proposed as the building block to model multilateral international negotiation processes [8].
The Friedkin-Johnsen (FJ) model has been validated in offline settings, where opinion evolution reflects direct interpersonal interactions and deliberations across various topics [9, 10]. However, online social environments introduce distinct mechanisms of opinion exchange: individuals are exposed to their contacts’ expressed opinions indirectly, primarily through public social media activities such as posts, comments, and shares [11, 12]. This mode of exposure frequently involves repeated encounters with viewpoints from a broader and more heterogeneous audience, which can heighten cognitive dissonance by presenting conflicting information [13, 14]. In online environments, opinion exposure is both asynchronous and large-scale, allowing conflicting signals to endure over time and reach diverse audiences. At the same time, empirical studies show that controversial or polarizing opinions propagate with exceptional speed and breadth compared to offline interactions, amplifying their influence on collective attitudes [15, 16]. These opinion expressions propagate through various content formats, including textual posts, news resharing, and multimedia, whose spread dynamics are effectively modeled by the theory of information cascades [17].
@@
Opinion dynamics and information cascades operate at distinct granularity and temporal scales. Information cascades focus on how individual pieces of information (e.g., a post, an image) are shared and propagate within the social network, while opinion dynamics concern the evolution of opinions through interactions among individuals. Both processes exhibit emergent behaviors, transitioning from local interactions to global effects such as content virality and polarization. In opinion dynamics, the specific social interactions leading to opinion change remain abstract and unspecified. Conversely, information cascades emphasize the sharing mechanism of content without addressing the effects of such sharing. Importantly, reposting online content received from friends (e.g., retweets) is one of the most common forms of interaction in digital platforms. Empirical evidence shows that in a complete 24-hour snapshot of Twitter activity, only 20.8% of tweets were original, while 50.7% were retweets, underscoring that resharing dominates user behavior [18]. This widespread resharing suggests a natural intertwining between opinion dynamics and information cascades: as users repeatedly encounter and propagate content, microscopic acts of resharing can influence macroscopic opinion evolution. Despite this evident overlap, the two processes have thus far been treated separately in the literature. In this work, we set out to fill this gap and establish an integrated model whereby the microscopic act of content sharing bears an effect in the macroscopic phenomenon of opinion evolution in online social networks.
To address this issue, we propose the Friedkin-Johnsen on Cascade (FJC), a novel model that integrates the FJ opinion dynamics framework with information cascades. Specifically, we hypothesize that opinions on a selected topic are expressed and disseminated through the same flow that governs the resharing of online posts. A post is initially shared by certain sources and then propagates as a cascade through successive neighborhoods until it loses relevance or reaches saturation. When a node reshapes a post (e.g., via a retweet with a comment or framing), it not only amplifies the content but also attaches its own expressed opinion on the topic—such as a textual remark, endorsement, or critique—thereby exposing neighbors to both the content and the sharer’s stance. Receiving such a reposted message from a neighbor can trigger an opinion update in the recipient, linking the micro-level cascade of reposting to the macro-level evolution of opinions.
First, we validated our FJC model on a real dataset of information cascades from Twitter, demonstrating its strong predictive power for online opinion dynamics. Building upon our previous work [7], we then analyzed the impact of cascade mechanisms on polarization, evaluating polarization levels using opinion vectors that maximize polarization in the classical FJ model. To this end, we tested the model on both synthetic and real social networks. Our main findings can be summarized as follows:
  • Highly central nodes emerge as opinion leaders with disproportionate influence, capable of either intensifying or mitigating polarization. Notably, regardless of the extent of information spread, the sequential nature of cascades causes central nodes to encounter content early, limiting their exposure to diverse opinions and making them less susceptible to external influence.
  • The FJC dynamics weaken the influence of peripheral nodes—even when they are numerous, stubborn, and aligned—since they often fail to reach and sway the central nodes.
  • Information cascades can substantially shape polarization outcomes, either amplifying or reducing them, depending on how opinions spread through the network.
These findings underscore how sharing mechanisms on online platforms introduce structural and temporal asymmetries that reshape opinion evolution and amplify the role of network topology compared to the classical FJ model. Furthermore, they highlight how online dynamics can exacerbate polarization beyond what is typically observed in offline social contexts. Neglecting the role of diffusion dynamics can therefore lead to serious misjudgments in assessing both opinion formation and polarization. These insights highlight the critical role of online information diffusion in shaping collective attitudes and social interactions and offer guidance for communication strategies in business, public discourse, and policy-making.
@@ II Related Works
The modeling of opinion dynamics has garnered the interest of researchers for decades, with this interest further amplified with the emergence of online social networks. There are several classes of opinion dynamics models, and the literature on the topic is extensive [19]. Here, we concentrate on the class of averaging models, which posit that users adjust their opinions based on those of the people around them to mitigate social discomfort. Among these models, the FJ model [3] is one of the most studied [20, 21, 22, 23, 24, 25, 26, 10, 27, 7]. Its inner workings are explained in Section III-C. Recently, research on the FJ model has taken two main directions. The first involves optimization problems based on FJ [28, 29, 30, 23, 20, 21], whose main goal is to make interventions into the social network to achieve specific desired outcomes in the final opinions. For example, some investigate how to modify the initial opinion of a set of nodes to maximise the opinion discord [28] or minimize the overall opinion [29]. The second direction explores the possible extensions of the FJ model to capture diverse observed features in the online social networks [31, 32, 33, 34, 35]. In particular, these papers incorporate both attractive and repulsive relationships between users [32, 34], even assuming that the influence of a node on another can vary according to the homophily principle [35]. Furthermore, Bhalla et al. in [31] analyzes the polarization effect of FJ on time-varying networks and Chitra et al. in [33] extended the FJ to reproduce filter bubbles. Differently from these works, we investigate the effect of the classical FJ model on a static network when the social debate is shaped by the temporal flow of news posts on online social platforms.
@@
To model the posting flow, we mainly take hints from the very popular independent-cascade model in [36] (see Section III-B for a description of how it works), which is able to reproduce realistic diffusion dynamics [37]. This seminal work has inspired many following studies and extensions [38, 39, 40, 41], which extend the classical model to include collaborative and/or antagonistic relationships among users, as well as dynamic networks. The main goal of information diffusion is to study the rate of diffusion of viral content depending on the parameters. Instead, we are not interested in the popularity of content, but rather in its diffusion process. Hence, we used the original independent-cascade model and analyzed its impact on the opinion formation process. Similarly to opinion dynamics, the literature on information diffusion is extensive, so we refer the interested reader to dedicated surveys for more details [42].
The works most closely related to ours are [43, 44], which also investigate the interplay between information diffusion and opinion formation. While [43] adopts the Friedkin–Johnsen (FJ) model and [44] employs a bounded confidence framework, both approaches differ fundamentally from ours. In [43], content influences opinions only once, by shifting initial biases based on content popularity; subsequent opinion dynamics unfold independently of further information diffusion. In [44], opinion adoption is driven solely by similarity to the root node’s opinion and is primarily used to model content virality, not the cognitive process of opinion change. In contrast, our FJC model assumes that every act of content sharing conveys the sharer’s current opinion, which in turn directly influences the recipient’s belief. This enables us to explicitly capture how the temporal structure of cascades shapes long-term opinion dynamics and polarization—an aspect not addressed in prior work.
@@ III-A The social network
We can describe an online social network platform as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the $v$ nodes $\mathcal{V}$ represent users and the edges $\mathcal{E}$ represent social connections between them. In the most general form, the graph $G$ is directed and the edges $(i,j)$ represent a follow relationship. Specifically, $(i,j)\in\mathcal{E}$ means that $i$ follows $j$, thus $j$ influences $i$ as its activity appears in $i$’s news feed. The set of nodes that has an influence on $i$ is thus composed of $i$’s out-neighbours, which we will denote with $\mathcal{N}^{+}(i)$. Instead, the set of nodes that is influenced by node $j$ is the in-neighbourhood of $j$, which we indicate with $\mathcal{N}^{-}(j)$ [^1] Please note that, generally, the neighbourhood of a node corresponds to our out-neighbourhood.. It is worth noting that in the special case of an undirected node, $\mathcal{N}^{+}(i)$ and $\mathcal{N}^{-}(i)$ are the same. This social connection can arise, for example, when $i$ follows $j$ on platforms like X or Facebook or when they are friends on Facebook. The strength of this relationship is mathematically represented by the quantity $\hat{w}_{ij}$, which is the $(i,j)$-th entry of the $n\times n$ matrix $\hat{W}$ that we call social matrix. The weight that node $i$ gives to its out-neighbours represents the degree of influence of their opinions. It is provided by the $(i,j)$-th entry of the $n\times n$ matrix $W$, named the influence matrix $W$, which is row-stochastic. On the $i$-th row, for example, we can see how the influence on $i$’s opinion is distributed among all nodes in $i$’s network. In situations where additional data is lacking, the matrix $W$ can be derived from $\hat{W}$ through row-normalization, expressed as $w_{ij}=\frac{\hat{w}{ij}}{\sum_{j}\hat{w}{ij}}$. This process capitalizes on the insight that stronger relationships should carry greater weight in opinion formation. It’s worth noting that, with this mechanism, even in cases where the social graph exhibits symmetry, the influence matrix results asymmetric.
@@ III-B Information cascade
Our model conceptualizes the act of posting social content following the independent cascade model [36]. In the independent cascade model, each node can be active or inactive. An initial set of active nodes $v$ has a single chance to activate their neighbours $w$, and each activation is successful with probability $p_{vw}$. The nodes $w$ that are successfully activated become active at the next step and try to activate their neighbours as well. The process continues until no other nodes can be activated. This framework is well-suited to modeling the propagation of shared elements in online social networks, such as tweets or posts, where each user typically has one opportunity to reshare content upon seeing it, and further reshares depend on their direct followers.
@@ III-C Friedkin-Johnsen opinion dynamic model
The FJ opinion dynamic model is a mathematical model that describes how people’s opinions change over time due to their mutual influences. It does this by representing people’s opinions as real numbers and then modelling two opposing forces that affect how those opinions evolve. The first force is an initial bias that each person $i$ has, which is represented by $u_{i}$. The second force is a desire to reduce the discomfort people feel when their opinions differ from those around them, as described by the cognitive dissonance theory. Each person $i$ has a parameter $\lambda_{i}$, called susceptibility, that determines how much weight they give to each of these forces. If $\lambda_{i}$ is 0, it means the person is completely stubborn and will never change their mind. If $\lambda_{i}$ is 1, it means the person is completely naive and will totally get over their initial bias. Node susceptibilities are generally represented by the diagonal matrix $\Lambda$, whose $i$-th diagonal element is equal to the susceptibility $\lambda_{i}$.
The FJ model is mathematically expressed as a simple update equation that is applied to each person’s opinion at discrete times. As discussed in our previous work [7], in the literature, the term “FJ model” refers to many variants of the original model. The update expression for the most general variant of the FJ model is given by the following equation:
@@
while the rest of the nodes maintain their opinion. Note that the term $(1-\gamma_{ij})x_{i}(t)$ balances the current expressed opinion of node $i$ against that of its neighbor $j$, with $\gamma_{ij}$ controlling the strength of influence from $j$. Here matrices $H=\{h_{ij}\}$ and $\Gamma=\{\gamma_{ij}\}$ replace $\Lambda$ and $W$ of the synchronous model, and are different from them. In fact, to guarantee the convergence of the model in Equation (3) to that in Equation (1), the formulation has to take into account the probability of node $i$ being randomly selected, which, in turn, depends on the degree of nodes. Since high-degree nodes are involved in more edges, when selecting edges at random high-degree nodes have a higher probability of being selected. For this reason, $H$ and $\Gamma$ should be defined from $\Lambda$ and $W$ as follows: